OpenStudy (strawberry17):

Find all angles(theta) between 1 and 2π for which cot(theta) = √3. How would you do this without a unit circle?

OpenStudy (tkhunny):

1) Why? 2) sqrt(3) = cot(x) = cos(x)/sin(x) = sqrt(3)/1 = [sqrt(3)/2]/[1/2] 3) If that doesn't look like a 30-60-90 right triangle, I don't know what does!

OpenStudy (strawberry17):

Where did you get the last values from? for #2 (the √3/2 ÷ 1/2)?

OpenStudy (tkhunny):

Magic? I happen to know that a 30-60-90 right triangle has sine and cosine equal to 1/3 and sqrt(3)/2, or the other way around. You should know this, too. Don't forget that this is in Quadrant I. There is another solution in Quadrant III where both values are negative?

OpenStudy (strawberry17):

Seems like I will need a lot of memorizing to do... Am I supposed to know automatically that this is in quadrant 1 from the positive value?

OpenStudy (tkhunny):

sine and cosine are positive. Absolutely you should know this is Quadrant I.

OpenStudy (strawberry17):

So what would be angle values?

OpenStudy (strawberry17):

It's still a bit confusing..

OpenStudy (tkhunny):

We established that cos(x) = sqrt(3)/2 and sin(x) = 1/2. That's 30º. You find the Quadrant III solution.

zepdrix (zepdrix):

Mmm ok these tangent and cotangent functions can be a little tricky :)

OpenStudy (strawberry17):

I agree! lol

zepdrix (zepdrix):

So like... let's think for just a sec. Lemme give a quick example. \[\Large \tan 60^o \quad=\quad \frac{\sin 60^o}{\cos60^o} \quad=\quad \frac{\left(\dfrac{\sqrt3}{2}\right)}{\left(\dfrac{1}{2}\right)}\quad=\quad \sqrt3\]

zepdrix (zepdrix):

If we had started with \(\Large \tan\theta=\sqrt3\) can you see how it would be tough to reference it back to sine and cosine since it's simplified down? That's what makes this problem a little tough :P heh

zepdrix (zepdrix):

If you remember your reference angles for tangent and cotangent, it makes this a little easier.

zepdrix (zepdrix):

\[\Large \cot \theta=\sqrt3\] Do you have this reference angle memorized or no? :D

OpenStudy (strawberry17):

no :/

zepdrix (zepdrix):

You and your strawberry ways -_-

OpenStudy (strawberry17):

hehe :D

zepdrix (zepdrix):

Hmmmmm how to explain this then :o hmmm

OpenStudy (strawberry17):

I wish I had them memorized...

OpenStudy (strawberry17):

You are very good at explaining... :D

zepdrix (zepdrix):

Try to memorize the reference angles for sine, cosine and tangent. Here are a couple important ones for tangent.\[\Large \tan \theta=\sqrt3 \qquad\to\qquad \theta=\pi/3\]\[\Large \tan \theta=\frac{1}{\sqrt3} \qquad\to\qquad \theta=\pi/6\]

zepdrix (zepdrix):

Since we are dealing with cotangent, we can relate it back to tangent using this identity:\[\Large \cot \theta =\frac{1}{\tan \theta}\]

zepdrix (zepdrix):

\[\Large \cot \theta=\sqrt3 \qquad\to\qquad \tan \theta=\frac{1}{\sqrt3}\] Understand what I did there? Since tangent is the `flip` of cotagent, I did a lil magic :o

OpenStudy (strawberry17):

ohhh, yess, I do see :D

OpenStudy (strawberry17):

you are so quick at typing up all that btw ~

zepdrix (zepdrix):

lol :3

zepdrix (zepdrix):

So as we listed above: \(\Large \tan \theta=\dfrac{1}{\sqrt3} \qquad\to\qquad \theta=\pi/6\)

zepdrix (zepdrix):

And this is the same as \(\Large \cot\theta=\sqrt3\).

zepdrix (zepdrix):

So our reference angle is \(\Large \theta=\pi/6\). Or 30 degrees, as you already figured out earlier it seems :)

zepdrix (zepdrix):

Hmm so we need the other one?

OpenStudy (strawberry17):

I think so, the one for quadrant 3??

zepdrix (zepdrix):

There are a couple of ways to approach this. We can use this,\[\Large \cot \theta \quad=\quad \frac{\cos \theta}{\sin \theta}\quad=\quad \sqrt3\] and then think about where else on the unit circle we're getting sqrt3. It will be where sine and cosine are both negative, right? That's how we can get a positive, when dividing negative by negative.

zepdrix (zepdrix):

There's another approach we can try though. This might be a tad easier.

zepdrix (zepdrix):

Since tangent and cotangent have period \(\Large \pi\), they will go through their full cycle of values in pi and then start to repeat after that. So we could simply add \(\Large \pi\) to find the other value.

zepdrix (zepdrix):

So our first angle was \(\Large \dfrac{\pi}{6}\), Our next one will be \(\Large \dfrac{\pi}{6}+\pi\)

zepdrix (zepdrix):

Too much maf? brain esplode? +_+

OpenStudy (strawberry17):

lol, yes way to much math for me that this hour... it's like 1:17am here :O but I need to turn in my quiz tomorrow and need to understand all this.. you are such a big help though ! :)

OpenStudy (strawberry17):

But I am not sure why it would be to add π to π/6..

zepdrix (zepdrix):

So think back to sine and cosine. Let's imagine we're at an angle of pi/6 on the unit circle. Try to picture it. If I tell you to add 2pi to that angle. I'm telling you to spin around the circle and last in the same spot, right?

OpenStudy (strawberry17):

yesss :)

zepdrix (zepdrix):

So if sin pi/6 = 1/2 What does sin (pi/6 + 2pi) =

zepdrix (zepdrix):

Hint: we landed in the same spot after we spun!! :O It should still give us 1/2 right?

OpenStudy (strawberry17):

ohhh, yea you are right

OpenStudy (strawberry17):

I was overthinking it...

zepdrix (zepdrix):

Well the same thing happens with the tangent/cotangent functions. But they start to repeat when you spin HALF WAY around the circle, not a full spin.

zepdrix (zepdrix):

cot (pi/6) = sqrt3 cot (pi/6 + pi) = sqrt3

OpenStudy (strawberry17):

Is that something I should memorize? :D

zepdrix (zepdrix):

The thing that I just wrote with the sqrt3? Or the thing about the half way spinning? :O

OpenStudy (strawberry17):

half way spinning :) lol

zepdrix (zepdrix):

Yes, try to remember it like this. Sine and Cosine have a period of 2pi. (They repeat their values after rotating 2pi more around the unit circle). Tangent and Cotangent have a period of pi. (They repeat their values after rotation pi more around the unit circle).

OpenStudy (strawberry17):

okay, I think this will really help me, yess! :) Thanks, it seems like the teacher does not explain as good as you :D

zepdrix (zepdrix):

Well the teacher can't give you personal 1 on 1 time ^^ So it can be hard to get every little detail hehe

OpenStudy (strawberry17):

sOo very true...

OpenStudy (strawberry17):

Was that all of the problem ??

OpenStudy (strawberry17):

I wish you could tutor me in for real.. I would be a pro in trig :D

zepdrix (zepdrix):

Ya I guess the way you would want to write your answer would be,\[\Large \theta=\frac{\pi}{6}, \quad\frac{7\pi}{6}\] Do you understand how I got the second angle? Did you have any trouble adding the pi to the fraction?

OpenStudy (strawberry17):

Yes fo sure I understand it with how well you explained it hehe. Thanks SO much!

zepdrix (zepdrix):

np \c:/ go to bed!!

OpenStudy (strawberry17):

:D